3. The demand for 1998 Honda Civics in BC is given by P = 5,400 – 200Q – 0.01I, where P is the car price in $, Q is quantity demanded (number of cars per year), and I is the average income of the buyers ($ per year). The supply is given by P = 1,000 + 200Q, where P is the car price in $, Q is quantity supplied (number of cars per year).

a. If the average income is I = $40,000 per year, what are the equilibrium price and quantity?

b. If the average income is I = $80,000 per year, what are the equilibrium price and quantity?

c. Is 1998 Honda Civic a normal or inferior good? How do you know?
1
Expert's answer
2019-10-01T12:31:08-0400

a. If I = 40,000, than demand is 

P = 5,400 - 200Q - 400.

At the same time supply is

P = 1,000 + 200Q.

From that equation we find:

4,000 - 200Q = 200Q


Q = 10 and P = 3,000


b. If I = 80,000, than demand is

P = 5,400 - 200Q- 800.

At the same time supply is 

P = 1,000 + 200Q.

From that equation we find:

3,600 - 200Q = 200Q


Q = 9, P = 2,800


c. 1998 Honda Civics is an inferior good, as with an increase of income, the quantity demanded of this good is decreasing.